how do you figure out a denominator?
11/24 + 5/? = 13/12

Answer by DrBeeee(684)

Replace your ? with x and get
(1) 11/24 + 5/x = 13/12 or
(2) 5/x = 13/12 - 11/24 or
(3) 5/x = (26 - 11)/24 or
(4) 5/x = 15/24.
Now cross multiply to get
(5) 24*5 = 15*x or
(6) x = 24*5/15 or
(7) x = 8
Let's check this with (1).
Is (11/24 + 5/8 = 13/12)?
Is (11/24 + 15/24 = 13/12)?
Is (26/24 = 13/12)?
Is (13/12 = 13/12)? Yes
Answer: ? = 8